OpenStudy (anonymous):

Multiply. (√(5)+4√(6))(5-4√(6))

OpenStudy (anonymous):

-57.7387676 ?

OpenStudy (saifoo.khan):

i was having a doubt.

OpenStudy (anonymous):

-91

OpenStudy (saifoo.khan):

still.

OpenStudy (anonymous):

I think it has to be the exact answers and u can use a radicals too

OpenStudy (anonymous):

you are right saifoo, something is wrong

OpenStudy (saifoo.khan):

hmm, it is.

OpenStudy (anonymous):

it's a differnence of tqo squared

OpenStudy (anonymous):

take it to its original

OpenStudy (anonymous):

I got 5√5-4√30+20√6-96... idk if it right

OpenStudy (anonymous):

maybe \[-96+5 \sqrt{5}+20 \sqrt{6}-4 \sqrt{30}\] you have it exactly

OpenStudy (anonymous):

so mines would be the correct answer?

OpenStudy (saifoo.khan):

yes.

jimthompson5910 (jim_thompson5910):

Is it really \[\large \left( \sqrt {5}+4\,\sqrt {6} \right) \left( \sqrt {5}-4\,\sqrt {6} \right)\] ???

OpenStudy (anonymous):

yes your answer is correct! good work

OpenStudy (anonymous):

there is a mistyping then

OpenStudy (anonymous):

for the given data the answer will be -57.7387676